26. (8 points) Consider a monopolist who faces a constant average and marginal cost of $5 and a linear demand curve of P = 20-20, where P is the price the monopolist charges and Q is the quantity consumers purchase. To obtain the optimal quantity and price, the monopolist needs to obtain the marginal revenue, which has the same intercept as the demand curve but twice as steep. Obtain the monopolists marginal revenue, optimal output, and price.
Answer #1

Consider the given problem here the demand curve is given by “P=20-Q”, => TR=P*Q.

=> TR = (20 - 2*Q)*Q = 20*Q - 2*Q^2, => MR = d(TR)/dQ = 20 - 4*Q. So, here we can see that “MR” have same intercept but have slope twice as steep as the demand. Now, there is a constant “MC” and “AC”, => MC = AC = $5.

=> At the equilibrium “MR” must be equal to “MC”, => MR=MC, => 20 - 4*Q = 5, => Q=(20-5)/4. So, the optimum quantity choice is given by “Q = 15/4 = 3.75”. So, the optimum price is given by, “P=20-2*Q = 12.5”. So, the optimum “P” and “Q” are given by “P=12.5” and “Q=3.75”.